OpenStudy (anonymous):

the average of two numbers,a and b is given by:a=a+b/2 if the average is 95 and number is 88 what is the other number? please show all of your work.

OpenStudy (asnaseer):

your question is slightly misleading, the average is given by:\[average=\frac{a+b}{2}\]

OpenStudy (asnaseer):

you have stated it as: a = a+b/2

OpenStudy (asnaseer):

the "a" on the left of the equation is not the same as the "a" on the right hand side

OpenStudy (anonymous):

(a+b)/2=95 Lets say a=88 (88+b)/2=95 Solve for b Multiply both sides by 2 88+b=190 Substract both sides 88 b=102 That is your other number.

OpenStudy (asnaseer):

so you just need to rearrange the equation to find "a" or "b" as follows:\[a+b=2*average\]\[a=(2*average)-b\]put in the values you are given to find "a".

OpenStudy (anonymous):

can i get some more HELP

OpenStudy (anonymous):

I already answere the question. You got that A=(a+b)/2 A=95 Any of a or b can be 88 So lets make a=88 Replacing 95=(88+b)/2 Multiply both sides by 2 190=88+b Substract both sides 88 102=b Thats your answer.

OpenStudy (anonymous):

WHERE DID YOU GET 95 FROM AND Y DID REPLACE THE A ON THE OUT OF THE EQUAL SIGN IT JUST NOT LOOKIN RIGHT THE ON THE OUTDSIDE NEVER CHANGE

OpenStudy (anonymous):

A=A+B/2

OpenStudy (anonymous):

WE SOLVING FOR A +B NOTHING ELSE

OpenStudy (anonymous):

A=Average a=firts number b=second number. Your question says. if the average is 95 and number is 88 what is the other number? Son replacing. A=95 a=88 b=?

OpenStudy (anonymous):

IM SORRY THE AVERAGE IS 93 I MESS THAT UP

OpenStudy (anonymous):

You got that A=(a+b)/2 A=93 Any of a or b can be 88 So lets make a=88 Replacing 93=(88+b)/2 Multiply both sides by 2 186=88+b Substract both sides 88 98=b Happy?

OpenStudy (anonymous):

HAPPY THANKS